# Understanding Trends, Patterns, and Dynamics in Global Acquisitions: A Network Perspective

### v2

@@ 4.8 Cross-city acquisition network
In this part, we construct the cross-city acquisition network, where nodes represent cities of organizations, and directed edges connect city A to city B if an organization from city A has acquired an organization in city B. The weight of each edge represents the number of acquisitions between the two cities. Figure 16 displays a sample of the cross-city acquisition network, including 50 cities with the highest numbers of connections (indegree + outdegree) with other cities.
In the visualization, node colors indicate their respective Louvain communities, and edge thickness is proportional to their weights. Additionally, node sizes correspond to their betweenness centrality, and the label sizes are proportional to their closeness centrality. According to our analysis, the network reveals three distinct communities: the green community represents cities in the United States, the golden community encompasses cities in California, and the pink community predominantly consists of Canadian and European cities, with a few Asian cities like Tokyo, Singapore, and Mumbai.
Figure 16: A sample of the cross-city acquisition network. Nodes are colored based on their community detected via the Louvain algorithm.
4.9 Centrality metrics
In this part, we calculate and discuss the centrality measures for the top ten central cities in the cross-city acquisition network. Figure 17 provides a detailed summary of centrality values for these cities. In terms of weighted degree centrality, New York exhibits the highest values, aligning with previous studies that assert New York’s predominant position in the global economy [57]. Following New York, London and San Francisco have the second and third highest weighted degrees, respectively. This observation coincides with Budd and Whimster’s findings, highlighting London’s status as a global financial center [58]. Additionally, Ester’s work suggests that San Francisco, with its accelerators in Silicon Valley, plays a significant role in attracting startups [59].
@@ 4.9 Centrality metrics
Figure 17: Side-by-side bar plot of centrality measures for the top ten central cities within the cross-city acquisition network.
@@ 4.10 Cross-border acquisition network
In this part, we construct the cross-border acquisition network to explore country relationships. We establish a directed edge from country A to country B whenever an organization in A acquires an organization in B. This creates a weighted directed network where edge weights signify the number of acquisitions between countries. Figure 18 showcases a sample of this network, featuring the top 60 countries based on their node degrees. Node size reflects their authority centrality, while label size corresponds to their betweenness centrality. Furthermore, nodes are color-coded according to their Louvain community. According to the findings, the United States, United Kingdom, and Germany exhibit the highest authority centralities respectively, indicating their significant influence in the global economy. Additionally, they possess the third-highest betweenness centralities, confirming their pivotal role in facilitating acquisition transactions among other countries. These results align with prior research, reinforcing the dominant position of the United States in the global economy [60, 61]. Moreover, previous studies have underscored the substantial contributions of the United Kingdom and Germany in cross-border acquisitions [62, 63].
Figure 18: A sample of the cross-border acquisition network. Nodes are colored based on their community detected via the Louvain algorithm.
@@
Figure 19: A sample of the cross-border acquisition network on the world map.
@@ 5 Conclusion
In this study, we analyzed global acquisitions with a network approach to gain insights into trends, patterns, and dynamics in this realm. We constructed the acquisition network and its projections on acquirers and acquirees to evaluate interconnections between companies. Our findings revealed that these networks have low densities and comprise a great number of small communities. By calculating assortativity coefficients, we demonstrated a preference among organizations to acquire companies from the same country, industry, or age bracket as themselves. ERGMs coefficients indicated the significant impact of different company features, such as country, region, city, and category, on the formation of connections between organizations, while founding date did not prove to be the cause of connections. Additionally, we conducted a temporal analysis on the acquisition network, which indicated that the average clustering coefficient has increased over time, while the network has become sparser, comprising more communities. We computed different centrality metrics for the cross-city acquisition network and found that New York, London, and San Francisco are global economic hubs. Finally, the analysis of the cross-border acquisition network revealed that the United States, United Kingdom, and Germany are the dominant countries in international acquisitions.
Future work can utilize insights from this paper in designing models for predicting future acquisitions. Additionally, researchers may focus on a specific city or country based on the economic hubs detected in this study to explore domestic acquisitions. Another avenue of research for future work is to create and analyze the merger network and compare its features with those of the acquisition network.
@@
The data that support the findings of this study are available from Crunchbase but restrictions apply to the availability of these data, which were used under the Crunchbase research license for the current study, and so are not publicly available. Data are however available from the authors upon reasonable request and with the permission of Crunchbase.

### v3

@@ 4.8 Cross-city acquisition network
Figure 13 displays a sample of the cross-city acquisition network, including 50 cities with the highest numbers of connections (indegree + outdegree) with other cities. In the visualization, node colors indicate their respective Louvain communities, and edge thickness is proportional to their weights. Additionally, node sizes correspond to their betweenness centrality, and the label sizes are proportional to their closeness centrality. According to our analysis, the network reveals three distinct communities: the green community represents cities in the United States, the golden community encompasses cities in California, and the pink community predominantly consists of Canadian and European cities, with a few Asian cities like Tokyo, Singapore, and Mumbai.
Figure 13: A sample of the cross-city acquisition network. The size of each node and its label are proportional to its betweenness and closeness centralities, respectively. Additionally, nodes are colored based on their community detected via the Louvain algorithm. This figure illustrates that the three detected communities mostly consist of cities in the United States, Europe, and California, respectively.
4.9 Centrality analysis
Figure 14 provides a detailed summary of centrality measures for the top ten central cities in the cross-city acquisition network. In terms of weighted degree centrality, New York exhibits the highest values, aligning with previous studies that assert New York’s predominant position in the global economy [17]. Following New York, London and San Francisco have the second and third highest weighted degrees, respectively. This observation coincides with Budd and Whimster’s findings, highlighting London’s status as a global financial center [10]. Additionally, Ester’s work suggests that San Francisco, with its accelerators in Silicon Valley, plays a significant role in attracting startups [23].
@@ 4.9 Centrality analysis
Figure 14: Side-by-side bar plot of centrality measures for the top ten central cities within the cross-city acquisition network. The number in front of each bar indicates the ranking of the respective city for each centrality measure. This figure reveals that New York, London, and San Francisco are the most central cities across different centrality metrics.
@@ 4.10 Cross-border acquisition network
Figure 15 showcases a sample of the cross-border acquisition network, featuring the top 60 countries based on their node degrees. Node size reflects their authority centrality, while label size corresponds to their betweenness centrality. Furthermore, nodes are color-coded according to their Louvain community. According to the findings, the United States, United Kingdom, and Germany exhibit the highest authority centralities respectively, indicating their significant influence in the global economy. Additionally, they possess the third-highest betweenness centralities, confirming their pivotal role in facilitating acquisition transactions among other countries. These results align with prior research, reinforcing the dominant position of the United States in the global economy [46, 21]. Moreover, previous studies have underscored the substantial contributions of the United Kingdom and Germany in cross-border acquisitions [33, 4].
Figure 15: A sample of the cross-border acquisition network. The size of each node and its label are proportional to its authority and betweenness centralities, respectively. Additionally, nodes are colored based on their community detected via the Louvain algorithm. This figure highlights the prominence of the United States, United Kingdom, and Germany in global acquisitions.
@@
Figure 16: A sample of the cross-border acquisition network on the world map. The size of each node and its label are proportional to its degree and betweenness centrality, respectively. The thickness of the edges corresponds to the number of acquisitions between the two countries. This figure further confirms the pivotal roles of the United States and the United Kingdom in the global economy.
5 Discussion
Before discussing the implications and limitations of our study and suggesting avenues for future research, we now address the research questions previously mentioned.
RQ1: What features influence the interconnections between organizations within the acquisition network?
We found that country and category group are the most influential features in the formation of edges (acquisition transactions) between companies. Previous studies also underlined the significant role of geographic proximity and industry on company acquisitions [3, 65]; however, our findings revealed that the impact of these factors remains significant even when considering a combination of company features. The reason behind this observation could be that acquiring firms are more likely to acquire companies in the countries and industries they have more experience with, particularly their own country and industry [3, 29].
RQ2: How has the structure of the acquisition network evolved over time?
We showed that the acquisitions network has become sparser overall, but the number of clusters within the network has increased, and these clusters have become denser, leading to a more complex structure. These patterns may be due to the absence of monopolies in the global economy and the shift towards industry-based acquisition transactions, which has accompanied the rise in the number of industries [30, 60, 44].
RQ3: Which cities or countries serve as global economic hubs in terms of acquisitions?
Our results revealed that New York, London, and San Francisco are the most central cities in global acquisitions, while the United States, United Kingdom, and Germany are the most influential countries in these transactions. This could be due to the fact that these cities and countries have open access policies, advanced infrastructure, and proximity to leading financial markets like Wall Street and Silicon Valley [17, 10, 23, 46, 33, 4].
5.1 Implications
The findings of our study provide valuable insights for policymakers aiming to create policies that promote global economic growth and success. Since most acquisition transactions occur within the same industries, policymakers can establish industry-specific regulations to enhance the growth and competitiveness of key sectors. Acknowledging New York, London, and San Francisco as economic hubs, policymakers in other cities can implement strategies modeled after these successful cities to attract acquisitions and investments. Additionally, considering the leading roles of the United States, United Kingdom, and Germany in international acquisitions, policymakers can ease cross-border transactions by lowering regulatory hurdles and encouraging other countries to engage in global acquisitions.
In addition to policymakers, acquiring firms and startups can also benefit from our results when making more informed acquisition decisions. For more successful integrations, acquirers should consider startups from the same country, industry, or age range. These firms can position themselves in economic hubs like New York, London, and San Francisco to gain access to more acquisition opportunities, partners, and investors. They should also consider expanding into key markets such as the United States, United Kingdom, and Germany, which can lead to higher success rates.
5.2 Limitations
We acknowledge that our study comes with certain limitations. The Crunchbase data we utilized has its own limitations and biases. There may be delays between the completion of venture capital deals and their disclosure. Some companies may not report the amount of investor capital lost, and startups that fail early might not create a profile on Crunchbase. Additionally, small local businesses and sole proprietorships might be underrepresented since Crunchbase tends to focus on high-growth technology and life sciences companies. Furthermore, as a US-based company, Crunchbase may have a bias towards companies from the US and other Western countries, resulting in less comprehensive data for companies from other regions [^5] https://news.crunchbase.com/methodology/. Therefore, the trends and patterns in acquisitions we observed in this study might be more accurate for Western companies and less generalizable to companies from underrepresented regions.
5.3 Future work
Future research can adopt a finer-grained perspective and focus on underrepresented countries to determine whether acquisition patterns in those regions differ from global trends. Another direction for future research is to complement Crunchbase data with local venture capital databases to ensure that local businesses and sole proprietorships are included in the analyses. Additionally, the insights from our study could be utilized in designing models for predicting future acquisitions.
@@ 6 Conclusion
In this study, we analyzed global acquisitions with a network approach to gain insights into trends, patterns, and dynamics in this realm. We constructed the acquisition network and its projections on acquirers and acquirees to evaluate interconnections between companies. Our findings revealed that these networks have low densities and comprise a great number of small connected components. By calculating assortativity coefficients, we demonstrated a preference among organizations to acquire companies from the same country, industry, or age bracket as themselves. ERGMs coefficients indicated the significant impact of different company features, such as country, region, city, and category, on the formation of connections between organizations, while founding date did not prove to be the cause of connections. Additionally, we conducted a temporal analysis on the acquisition network, which indicated that the average clustering coefficient has increased over time, while the network has become sparser, comprising more weakly connected components. We computed different centrality metrics for the cross-city acquisition network and found that New York, London, and San Francisco are global economic hubs. Finally, the analysis of the cross-border acquisition network revealed that the United States, United Kingdom, and Germany are the dominant countries in international acquisitions.
Abbreviations
IT, Information Technology; SNA, Social Network Analysis; ERGM, Exponential Random Graph Model; IPO, Initial Public Offering; AUC, Area under the ROC Curve; AI, Artificial Intelligence; CV, Curriculum Vitae; M&A, Mergers and Acquisitions; NLP, Natural Language Processing; AIC, Akaike Information Criterion.
@@
The data that support the findings of this study are available from Crunchbase but restrictions apply to the availability of these data, which were used under the Crunchbase research license for the current study, and so are not publicly available. However, the analysis scripts supporting the conclusions of this paper are available at https://github.com/kalhorghazal/Company-Acquisition-Paper.
@@
Acknowledgements
The authors express gratitude to Crunchbase for facilitating access to data.